OpenStudy (anonymous):

Suzie's budget is: hop+on+the+bus=gus Knowing that each letter stands for a different digit 0 through 9 and no two letters stand for the same digit. How much money did Suzie request?

OpenStudy (anonymous):

ok im working on it... heres what i got so far, hop+on+the=gus-bus ( us -us) has to equal the letters for zero, whitch also means that op+on+he = 00

OpenStudy (anonymous):

thanks for working on it... I trying to come up with an answer based on your tips so far

OpenStudy (anonymous):

i shouold say op+on+he = a number that has zero in the digit and tens place

OpenStudy (anonymous):

[(h)*10^2]+[(o)*10^1]+[(p)*10^0]+[(o)*10^1]+[(n)*10^0]+[(t)*10^2]+[(h)*10^1]+[(e)*10^0]=(g-b)x10^2

OpenStudy (anonymous):

[(h)*100]+2[(o)*10]+p+n+[(t)*100]+[(h)*10]+e=(g-b)x100

OpenStudy (anonymous):

reading your reply...

OpenStudy (anonymous):

Sorry... what number does each letter represent?

OpenStudy (anonymous):

im not sure, what i did is assume that xyz as a number in this system would correspond to x = the hundreds value, y = tens, and z = digit, so i sorted them out like that, now im going to try to divide it out

OpenStudy (anonymous):

110h+20o+p+n+100t+e=100g-100b, this is proabally as far as i can get lol im gonna keep trying

OpenStudy (anonymous):

what class is this for:? linear algebra?

OpenStudy (anonymous):

Algebra 1 in 7th grade

OpenStudy (anonymous):

working through your reply now...

OpenStudy (anonymous):

7th grade algebra!?

OpenStudy (anonymous):

i think my methods are too abstract!, im at a loss, im sorry

OpenStudy (anonymous):

yes... but this is a high school credit even though I am in 7th grade. crazy, I know.

OpenStudy (anonymous):

actually, you are helping me think about it in a different way... thanks.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@phi

jimthompson5910 (jim_thompson5910):

This is just something where you have to use trial and error notice that p+n+e+s = xs where the x is the carry The carry will be larger than 0 because you're starting at some number s and you're looping back to s. The only way to do this is if the carry is at least 1.

jimthompson5910 (jim_thompson5910):

oh and because you're dealing with single digit numbers, x = 1 because there's no possible way to have a carry larger than 1

jimthompson5910 (jim_thompson5910):

this would then mean 1+o+n+h+u = yu where y is the new carry, again because we start from u and end back at u, we must have carry, which is also 1

jimthompson5910 (jim_thompson5910):

finally one last thing to realize is the sum is a 3 digit number, so the sum of the hundreds digits must not exceed 10 (keep the carry in mind)

OpenStudy (anonymous):

thanks @jim_thompson5910